### v4

# When Algorithms Meet Artists: Semantic Compression of Artists’ Concerns in the Public AI-Art Debate

Abstract
[…]
1 Introduction
1.1 The Paradoxical Position of Artists
When new technologies emerge, whose voices shape public understanding? This question is central to technology governance: if public discourse systematically underweights certain stakeholder groups, collective understanding becomes misaligned with stakeholder reality—a form of epistemic failure with policy consequences. We investigate this question in a consequential domain: generative AI and creative labor.
Artists occupy a paradoxical position in this technological transition. They stand as both the primary authors of training data for these systems and the stakeholders most susceptible to the systems’ disruptive impacts. Generative models are fundamentally reshaping the production, circulation, valuation, and attribution of creative work, while also intensifying persistent disputes concerning consent, compensation, authorship, and the legitimacy of data practices (Lovato et al., 2024; AOI Coalition of Artists, 2025; Creators Coalition on AI, 2025). The effects of these disputes have become increasingly tangible. In 2023, a landmark class-action lawsuit (Andersen v. Stability AI) brought artists’ grievances before a federal court, alleging that generative image models developed using copyrighted works without consent constitute mass infringement (Andersen et al. v. Stability AI et al., 2023). As Varvasovszky and Brugha (2000) emphasize, identifying and understanding primary stakeholders is essential for evaluating the social consequences of technological change. For artists, this transformation raises fundamental questions about creative ownership, labor, and artistic identity (McCormack et al., 2019; Browne, 2022).
1.2 Stakeholder Frames: The “Territory” of Artist Concerns
Relevant research shows that we cannot treat “artists” as a single, unified group in their interpretation of AI developments. Studies by Lovato et al. (2024) and Kawakami and Venkatagiri (2024) reveal that artist outlooks are actually highly specific and often internally complex. Lovato et al. (2024) operationalize these concerns across five dimensions (Utility, Ownership, Compensation, Transparency, and Threat) comprising 34 distinct “frames,” understood here in Entman’s (1993) sense as interpretive packages that select and emphasize particular aspects of a perceived reality. This granularity is essential: one artist may embrace AI tools while rejecting the economic models of companies that develop them; another might support transparency requirements but remain uncertain about ownership claims. Similarly, Kawakami and Venkatagiri (2024) document the spectrum of impacts that generative AI has on creative practitioners, ranging from the benefits of workflow augmentation to the considerable employment risks that accompany its widespread implementation, while Jiang et al. (2023) identify distinct patterns of concerns across different artistic communities.
These dimensions matter not only because they document disagreement among artists, but also because they provide the empirical basis to define stakeholder frames—coherent interpretive positions that can be systematically compared with the accounts that dominate broader public discourse.
1.3 Public Discourse: The “Map” that Compresses the Territory
However, the “map” of public discourse often fails to capture this territory. Public understanding of new technologies does not emerge neutrally: it is shaped by media agendas, institutional power, and visibility regimes that determine which issues become salient and which voices appear legitimate. Public discourse functions as a form of distributed cognition, aggregating information and shaping collective beliefs about new technologies. When this aggregation systematically underweights certain stakeholder groups, collective understanding becomes misaligned with stakeholder reality.
Zai et al. (2025) examined AI debates across Western media, finding that economic and scientific actors possess disproportionate standing, systematically foregrounding narratives of “progress” and economic potential while overshadowing ethical and social frames. Similarly, Bøgh (2025) found that news media tend to reduce the complex AI-art interface into binary oppositions—“AI as Threat” versus “AI as Tool”—stripping away the nuanced labor concerns held by practitioners. Banks and Li (2025) extend this analysis by mapping public debates about image-generative AI specifically, revealing how particular framings achieve dominance while others remain peripheral.
The structural mechanics of the contemporary public sphere exacerbates this phenomenon. Sichach (2025) argues that we have transitioned from traditional editorial gatekeeping to “algorithmic agenda-setting,” where recommender systems prioritize content that maximizes engagement rather than epistemic completeness. Pane (2025) describes this as a “post-mediatized” sphere in which algorithmic curation establishes the epistemic boundaries of discourse, determining which voices are amplified and which remain invisible. Huang and Gadavanij (2025) demonstrate that in AI-related media coverage, not all social actors are represented equally, and limited coverage contributes to systematic underrepresentation of affected stakeholders. Together, the evidence suggests that the social and structural conditions exist for the systematic marginalization of artist voices—a hypothesis that has not been empirically tested in the AI-art domain using computational methods that can quantify the degree and nature of this marginalization.
1.4 Epistemic Injustice: Marginalization Without Explicit Exclusion
Recent scholarship on epistemic injustice in AI governance provides a useful theoretical lens. Baeyaert (2025) demonstrates that when knowledge produced by affected groups is systematically backgrounded relative to institutional or technical narratives, marginalization can occur even without explicit exclusion—through differential amplification, topic selection, and standing asymmetries rather than overt silencing. In the AI-art debate, where artists are directly affected by data practices and market restructuring, epistemic marginalization is a plausible outcome of the visibility regimes documented by Sichach (2025) and Pane (2025). The question is not whether artists are formally prohibited from speaking, but whether their concerns achieve adequate salience within the discursive ecosystem that shapes public understanding and policy attention.
1.5 Temporal Context: From Novelty to Crisis
The timing of this question matters because AI-art discourse has shifted sharply over the last decade. Early public discussion (2013–2020) disproportionately emphasized novelty, experimentation, and philosophical speculation about machine creativity, framed by milestones such as Google’s DeepDream (Mordvintsev et al., 2015) and the emergence of Generative Adversarial Networks (Goodfellow et al., 2014), which spawned discussions of algorithmic aesthetics and machine consciousness. This period generated substantial scholarly attention to questions of algorithmic aesthetics, machine consciousness, and the ontological status of AI-generated works (Browne, 2022; Mazzone and Elgammal, 2019). Mazzone and Elgammal (2019) explored the creative potential of AI systems, while Browne (2022) interrogated the very category of “AI artist.” The 2018 Christie’s auction of Portrait of Edmond de Belamy—a GAN-generated image—brought AI art into mainstream cultural discourse, intensifying debates about authorship and market legitimacy.
The more recent period (2021–2025), characterized by the mass adoption of text-to-image systems like DALL-E (Ramesh et al., 2021) and latent diffusion models (Rombach et al., 2022), has seen a pivot toward labor disruption, legal conflict, and institutional integration. Cheung (2024) shows that media environments frequently oscillate between “opportunity” and “crisis” narratives, an alternation that can simplify stakeholder realities into a small set of repeatable tropes. Saeedi and Taleghani (2025) apply topic modeling to examine the discourses surrounding generative art diffusion, revealing how particular themes achieve prominence at different temporal moments. However, their analysis focuses on the internal structure of public discourse without systematically comparing it to empirically-derived stakeholder positions. Despite this shift, we lack systematic empirical evidence on the alignment between public narratives and stakeholder priorities: does public discourse reflect the distribution of artists’ concerns, or does it compress them into a narrow semantic region?
1.6 Research Gap: Semantic Compression as Measurable Marginalization
This gap is especially important for computational social science because it connects a normative claim—that artists are primary stakeholders whose voices should inform AI governance—to a measurable outcome: representation and salience within public discourse. This gap also connects to broader questions in behavioral science about how collective attention is allocated and how stakeholder knowledge is aggregated (or fails to aggregate) in public deliberation. Existing research often focuses on technological capabilities, philosophical questions of machine creativity, or legal analysis of copyright (Saeedi and Taleghani, 2025; Banks and Li, 2025; Kawakami and Venkatagiri, 2024; Mazzone and Elgammal, 2019; Jiang et al., 2023), but seldom measures how stakeholder concerns are geometrically distributed within the broader discursive ecosystem. As DiMaggio et al. (2013) argue, topic modeling offers a powerful tool for exploiting affinities between computational text analysis and the sociological study of culture, enabling researchers to map meaning structures at scale while remaining attentive to the relational positions of different actors within those structures.
We still lack systematic evidence on whether public attention tracks the distribution of stakeholder frames or instead produces what we term semantic compression: the collapse of a diverse set of artist concerns into a narrow region of the public meaning space. If public discourse acts as a “low-pass filter” on stakeholder complexity, amplifying broad tropes while attenuating nuanced positions, the consequences for AI governance are significant. Policymakers, platforms, and institutions that rely on public discourse as a proxy for stakeholder priorities would systematically underweight the concerns of affected communities.
1.7 This Study: Mapping Discourse, Projecting Stakeholders
In this paper, we address this gap by mapping the semantic landscape of public discourse on AI and art (2013–2025) and projecting artist survey statements into that space to quantify alignment, separation, and salience. Our public corpus spans 131 documents and 891 processed text chunks across news, podcasts, panel discussions, legal filings, and research papers—yielding 22 distinct discourse topics. We operationalize stakeholder frames using the five dimensions identified by Lovato et al. (2024), comprising 34 distinct frames from 252 US-based practicing artists.
Hypotheses
  1. 1.
    H1: Artist frames will concentrate in a small subset of public discourse topics, indicating semantic compression.
  2. 2.
    H2: Distributional divergence between artist frames and public discourse will persist after style controls, indicating semantic marginalization.
  3. 3.
    H3: Governance-related artist concerns (ownership, transparency, compensation) will show greater under-representation than affective concerns (threat, utility).
2 Methods
2.1 Study Design Overview
We measured whether artist stakeholder concerns achieve proportional representation in public discourse about generative AI and art. Our approach constructs a semantic reference map from public discourse (2013–2025) and projects artist survey responses into that shared space to quantify representational alignment—an approach aligned with recent computational methods for comparing how different stakeholder groups articulate contested topics (Elmholdt et al., 2025; Matsui and Ferrara, 2024).
We operationalize artist concerns using five dimensions from prior survey research (Lovato et al., 2024): Threat, Utility, Transparency, Ownership, and Compensation. These dimensions structure our extraction of artist probes, generation of style-matched comparison probes, and interpretation of theme-specific misalignment.
2.2 Data Sources and Corpus Construction
2.2.1 Public Discourse Corpus
We assembled a multimodal corpus covering 2013–2025 using targeted Google Search queries related to generative AI and art (e.g., “AI art impact on artists,” “generative AI copyright”; full query list in Supplementary Table S1). For each query, we sampled all unique documents appearing on the first page of results. Searches were conducted in a logged-out browser environment to reduce personalization artifacts.
We included English-language sources from news articles, blogs, podcasts, interviews, panel discussions, peer-reviewed papers, and legal filings. We excluded social-media-first platforms (Reddit, Twitter/X) to maintain comparability with search-ranked and editorialized sources (Sichach, 2025; Pane, 2025). Spoken sources were transcribed using OpenAI Whisper with manual review.
The final corpus comprised 131 documents segmented into 891 text chunks ($\sim$250 words each with 25-word overlap). Each chunk was annotated with publication year and media type for stratified robustness testing.
2.2.2 Artist Survey Corpus
Artist stakeholder language was derived from the survey dataset of Lovato et al. (2024), which captured practicing artists’ attitudes toward generative AI. We filtered to US-based respondents actively practicing as artists ($n=252$) to match the geographic scope of our public corpus.
To enable direct comparison in shared semantic space, we converted survey responses into short declarative probes using fixed templates (e.g., “I agree that AI art models are a threat to art workers”). This produced 1,259 artist probes across 34 distinct frame combinations spanning the five concern dimensions.
2.2.3 Public Probes (Style Control)
A key confound in comparing survey statements to public discourse is stylistic mismatch. To isolate semantic content from discourse format, we extracted style-matched public probes—short declarative sentences from public documents that match the syntactic form of artist probes.
Using GPT-5.1, we generated 250 synthetic Likert-style anchor statements following a factorial design: 5 themes $\times$ 5 agreement levels (strongly disagree to strongly agree) $\times$ 10 discourse-style variants (e.g., blog opinion, news editorial, artist interview, legal brief; see Supplementary Table S4 for design matrix and representative examples; full set available in our code repository). We then retrieved the nearest-neighbor sentences from the public corpus for each anchor. A human-LLM validation filter (Ziems et al., 2024) retained only substantively on-theme sentences, yielding 379 public probes. This enables direct comparison: if artist-public divergence persists when both are expressed as short declaratives, the mismatch is semantic rather than stylistic.
2.3 Semantic Mapping
All text units were encoded using the sentence-transformers/e5-large-v2 embedding model, selected for robust performance across diverse genres.
To ensure stable cross-corpus comparison, we implemented a consensus-based approach (Supplementary Figure S1): we generated 31 UMAP projections with different random seeds, computed pairwise distance matrices for each, averaged these matrices, and fit a final 8-dimensional embedding from the consensus distance structure (see Supplementary Methods for details and stability validation). This approach increases average seed-to-consensus Adjusted Rand Index from 0.56 (naive coordinate averaging) to 0.71 (distance-matrix consensus), yielding a more stable reference geometry into which new data can be projected without distorting the base map.
We trained a projection head (multi-layer perceptron) to map from embedding space to consensus coordinates, enabling placement of artist probes and public probes into the reference space (validation $R^{2}=0.73$; $k$-NN neighborhood preservation $82.4\%$).
The 891 public chunks were clustered using HDBSCAN, yielding 22 distinct topic clusters that form the semantic inventory for all comparisons.
2.3.1 Methodological Contribution
The consensus-based projection approach we introduce addresses a key challenge in cross-corpus semantic comparison: ensuring that stakeholder statements can be reliably placed into a reference discourse space without distorting the underlying geometry. Standard UMAP projections are sensitive to random seed initialization, producing unstable coordinates that complicate cross-study comparison. Our consensus methodology—averaging distance matrices across multiple seeded projections before fitting the final embedding—yields substantially more stable reference geometries (seed-to-consensus Adjusted Rand Index: 0.71 vs. 0.56 for naive coordinate averaging). The MLP projection head enables out-of-sample placement with strong neighborhood preservation (82.4% $k$-NN consistency). This methodology is currently being applied to stakeholder representation in other technology domains (platform labor, medical AI), and a dedicated methods paper describing the approach in detail is in preparation.
2.4 Topic Interpretation
Each cluster was interpreted independently by five annotators: two human coders (first author and research assistant) and three LLMs (GPT-5.1 Pro, Gemini Pro 1.5, Claude Opus 4.5). This multi-annotator approach addresses subjectivity in topic labeling (Ziems et al., 2024). A final judge (GPT-5.1 Pro) consolidated interpretations into final labels, keywords, and macro-thematic groupings (see Supplementary Table S2 for full topic inventory).
2.5 Inference Framework
We employed three complementary analyses to establish robustness:
2.5.1 Distributional Analysis
We compared how corpora distribute across the 22 topics using chi-square tests, Cramér’s $V$ (effect size), and Jensen-Shannon divergence. Pairwise comparisons tested: Public vs. Artist Probes (raw divergence), Public Probes vs. Artist Probes (style-controlled divergence).
2.5.2 Geometric Analysis
We computed centroid distances between corpora in 8D space and $k$-nearest-neighbor same-source rates. Permutation tests (10,000 iterations) generated null distributions for significance testing.
2.5.3 Robustness Checks
We tested whether divergence patterns could be explained by media type or publication year using within-stratum association tests.
2.6 Salience Ratios
For each topic, we computed salience ratios: the proportion of artist probes in that topic divided by the proportion of the comparison corpus. Ratios >1>1 indicate artist overconcentration (public under-emphasis); ratios <1<1 indicate public overconcentration.
To quantify theme-level misalignment, we identified the minimal cluster sets capturing 80%, 90%, and 95% of each artist theme (Threat, Utility, Ownership, Transparency, Compensation) and computed salience ratios for these theme-regions.
2.7 Use of Large Language Models
We used LLMs (GPT-5.1 Pro, Gemini Pro 1.5, Claude Opus 4.5) for anchor generation, validation filtering, and cluster interpretation. In all cases, human oversight was retained, and we employed multiple models to mitigate single-model biases (Ziems et al., 2024). Artist survey data are available through Lovato et al. (2024).
3 Results
3.1 Overview
We constructed a semantic map of public AI-art discourse from 131 documents (891 text chunks) spanning 2013–2025, including news articles, podcasts, panel discussions, academic papers, and legal filings. Clustering identified 22 distinct topics organized into five macro-thematic groups (Table 1). The largest group—Institutions and Markets (35.6%)—reflects the prominence of curatorial, commercial, and spectacular narratives in public visibility. Full topic descriptions are provided in Supplementary Table S2.
Table 1: Macro-thematic groupings of the 22 public discourse topics, with share of public corpus and number of constituent topics.
Macro-theme Share of corpus Topics
Institutions & Markets 35.6% 7
Governance & Rights 26.0% 5
Technical Genealogy 18.5% 5
Practice & Pedagogy 12.4% 3
Philosophy of Creativity 7.5% 2
Into this semantic space, we projected 1,259 probes derived from survey responses of 252 US-based practicing artists, capturing 34 distinct concern frames across five dimensions: Threat, Utility, Ownership, Transparency, and Compensation. We also projected 379 style-matched public probes—short declarative sentences extracted from public discourse—to control for stylistic differences between survey and media formats.
3.2 Artist Concerns Concentrate in a Narrow Discursive Region (H1)
Figure 1: Artist concerns concentrate in a narrow region of public discourse. Distribution of 1,259 artist probes across 22 public discourse topics. Just 4 topics capture 95.4% of artist concerns, while 14 topics (62.4% of public discourse volume) contain no artist perspective whatsoever.
When artist probes are projected into the public discourse semantic space, a stark concentration pattern emerges (Figure 1). Of the 22 public discourse topics, just 4 capture 95.4% of artist concerns:
  • Cluster 17 (Artist backlash and ethical gatekeeping): 40.6%
  • Cluster 10 (Pop-culture existential AI debate): 21.3%
  • Cluster 16 (AI as artistic collaborator): 18.7%
  • Cluster 9 (DALL-E “democratization” debate): 14.9%
The remaining 18 topics capture only 4.6% of artist probes. More strikingly, 14 topics contain zero artist perspective whatsoever—and these 14 topics comprise 62.4% of public discourse volume. The regions where artists are absent are dominated by institutional narratives (museum exhibitions, NFT markets, funding models), technical genealogies (DeepDream, GANs, diffusion models), and philosophical speculation about machine creativity.
This concentration is not an artifact of limited artist diversity. Artists in our sample articulate 34 distinct frames representing substantively different positions: some embrace AI as a creative tool while others view it as an existential threat; some would donate their work freely while others demand ongoing royalties; some assert exclusive ownership rights while others reject property frameworks entirely. Yet this heterogeneity—the “territory” of stakeholder reality—collapses into a narrow region of the public discourse “map” (Figure 2).
Figure 2: Key mismatches between public discourse and artist concerns across the macro-thematic groups. Public discourse is dominated by Institutions & Markets (35.6%) and Technical Genealogy (18.5%), while artist concerns concentrate heavily in Governance & Rights (57.5%).
The compression is particularly acute for governance-related concerns. All 18 ownership frame combinations—representing positions from “Artist Exclusive” to “Shared: User + Artist” to “Anti-Ownership/Commons”—map to a single topic (Cluster 17). The 10 compensation frames similarly concentrate in just 2–3 topics. Public discourse has captured the issue of AI art but stripped it of its internal political structure, reducing complex stakeholder positions to broad tropes of “backlash” or “democratization.”
These findings support H1: Artist frames concentrate in a small subset of public discourse topics, with a compression ratio of approximately 8.5:1 (34 frames $\rightarrow$ 4 primary topics).
3.3 Marginalization Is Semantic, Not Stylistic (H2)
A potential concern is that the observed separation reflects stylistic mismatch—survey responses are short declaratives, while public discourse includes long-form articles and transcribed speech—rather than substantive semantic differences. We addressed this by comparing artist probes to style-matched public probes: short declarative sentences extracted from public documents that match the syntactic form of survey responses.
If the artist–public separation were primarily stylistic, it should collapse when both corpora share the same format. It does not.
Distributional evidence.
Artist probes and public probes show large distributional divergence ($\chi^{2}$ $p<10^{-115}$; Cramer’s V = 0.606). While this is reduced from the raw public–artist comparison (Cramer’s V = 0.750), it remains well above the threshold for large effects (V >0.5>0.5). Jensen–Shannon divergence drops from 0.338 to 0.200—a meaningful reduction, but far from collapse.
Geometric evidence.
In 8-dimensional semantic space, artist probes and public probes remain separated (centroid distance = 1.336), and 76.5% of public probes have other public probes as their nearest neighbors rather than mixing with artist probes. When comparing artist probes to raw public discourse, 92.1% of artist probes have other artist probes as nearest neighbors—indicating that artists occupy a geometrically distinct region of the semantic manifold rather than merely expressing the same ideas differently.
Interpretation.
Even when public discourse is “chopped” into survey-style sentences, it does not overlap with what artists are saying. The public and artists are not discussing the same issues from different angles; they are discussing different things entirely. Institutional and technical narratives dominate public discourse even at the sentence level, while artist concerns about consent, compensation, and creative rights remain peripheral.
These findings support H2: Distributional and geometric divergence persist after style controls, confirming that marginalization is semantic rather than stylistic.
3.4 Governance Concerns Are Most Marginalized (H3)
Not all artist concerns are equally underrepresented. We computed salience ratios—the concentration of artist probes in theme-relevant regions divided by the concentration of public discourse in those same regions—to quantify differential marginalization across concern dimensions.
Table 2: Theme-based salience ratios at 90% coverage threshold. Values >1>1 indicate artist overconcentration relative to public discourse (i.e., public under-emphasis of artist concerns).
Theme SR (vs. Public) SR (vs. Public Probes)
[…]
The pattern is clear: The concerns most actionable for policy—transparency requirements, ownership claims, compensation mechanisms—are precisely those most systematically absent from public discourse. Affective reactions to AI (“it’s threatening” or “it’s useful”) achieve some public salience, but the specific governance demands that could inform regulation remain peripheral.
These findings support H3: Governance-related artist concerns (ownership, transparency) show greater underrepresentation than affective concerns (threat, utility). The most policy-relevant concerns are the most marginalized.
3.5 Summary
Three convergent lines of evidence establish that artist stakeholder concerns are systematically compressed in public AI discourse:
  1. 1.
    Concentration (H1): 95.4% of artist frames map to just 4 of 22 topics; 14 topics (62.4% of public discourse) contain no artist perspective.
  2. 2.
    Semantic basis (H2): Divergence persists after controlling for stylistic differences between survey and media formats (Cramer’s V = 0.606; centroid distance = 1.336).
  3. 3.
[…]
The 34 distinct frames articulated by artists—representing substantively different positions on threat, utility, transparency, ownership, and compensation—collapse into a narrow region of public meaning-space. In the Discussion, we interpret this pattern as a form of epistemic marginalization with implications for AI governance.
4 Discussion
4.1 Overview: What the Findings Imply
Our results provide strong evidence that public discourse on AI and art acts as a low-pass filter on stakeholder complexity. While artists articulate a high-dimensional landscape of 34 distinct frames spanning nuanced positions on threat, utility, ownership, transparency, and compensation (Lovato et al., 2024), public discourse compresses this variance into a narrow region of the semantic map. Across distributional, geometric, and stratified analyses, we find that artist stakeholder frames are not proportionally represented in the public discourse topic landscape, even after accounting for stylistic form.
This pattern is consistent with recent media research showing that public debates about AI grant disproportionate standing to economic and scientific actors while marginalizing ethical and social frames (Zai et al., 2025), and with arguments that agenda-setting is increasingly shaped by algorithmic curation and engagement-driven ranking rather than epistemic completeness (Sichach, 2025; Pane, 2025). Huang and Gadavanij (2025) document analogous dynamics in AI-in-education discourse, where affected stakeholders receive systematically less coverage than institutional actors despite the direct relevance of their concerns. The core empirical signature we document—95% of artist probes concentrated in 4 of 22 public topics—provides a measurable instantiation of what Baeyaert (2025) terms marginalization without explicit exclusion: affected-group knowledge is systematically backgrounded relative to institutional and technical narratives even when present in the discourse environment.
The magnitude of this compression is striking. Artists in our sample hold internally diverse and often conflicting positions: some embrace AI as a creative tool, while others view it as an existential threat; some would donate their work freely, while others demand ongoing royalties; some support mandatory transparency, while others remain uncertain about ownership claims (Lovato et al., 2024; Kawakami and Venkatagiri, 2024). Yet this heterogeneity—the “territory” of stakeholder reality—maps onto a small fraction of the public discourse “map.” The remaining 18 topics, comprising over 60% of public discourse volume, contain virtually no artist perspective.
4.2 The Phenomenon of Semantic Compression
4.2.1 Public Discourse Is Broad, But Salience Is Uneven
The public discourse landscape on AI art appears pluralistic at first. Our semantic map contains 22 topics spanning technical genealogies, institutional and market narratives, governance and law, philosophy of creativity, and tool practice. This breadth is consistent with recent topic-mapping work showing that public debates about image-generative AI contain multiple distinct interpretive clusters rather than a single linear narrative (Saeedi and Taleghani, 2025; Banks and Li, 2025).
However, the key question is not whether multiple topics exist, but which topics become salient and which frames gain standing—a distinction central to framing theory (Entman, 1993). Prior work suggests that AI discourse frequently foregrounds progress/economic potential and technologically oriented narratives while ethical-social concerns receive less attention or are simplified (Zai et al., 2025; Bøgh, 2025). Bøgh (2025) specifically documents this pattern in AI-art media coverage, finding that nuanced labor concerns are reduced to “AI as Threat” versus “AI as Tool” framings. Our results extend this logic by showing that stakeholder-derived frames occupy a comparatively narrow region of the public topic space: 4 topics account for 95% of artists’ concerns, while 14 topics contain none whatsoever.
4.2.2 Agenda Misalignment Manifests as Semantic Compression
When we project artist survey statements into the public semantic map, artist content concentrates sharply in Clusters 9 (DALL-E democratization debate), 10 (pop-culture existential AI art), 16 (AI as artistic collaborator), and 17 (artist backlash and ethical gatekeeping). We term this concentration semantic compression: a high-dimensional set of stakeholder positions collapses into a narrow set of public discourse regions.
Importantly, this is not a claim that artists lack nuance. Lovato et al. (2024), Kawakami and Venkatagiri (2024), and Jiang et al. (2023) document fine-grained and internally divergent positions. Instead, our contribution is to show that the public discourse “map” does not allocate proportional representational space to this stakeholder “territory.” In the public sphere, the distinction between an artist who would donate their work freely (the “Commons Donor” frame) and an artist who demands profit-sharing (the “Value Capture” frame) is erased; both land in Cluster 9. The 18 distinct ownership frames—ranging from exclusive artist rights to shared user-artist claims to rejection of all property frameworks—all collapse into Cluster 17. The public agenda has “captured” the issue of AI art but has stripped it of its internal political logic, reducing complex labor disputes to broad tropes of “resistance” or “democratization.”
4.2.3 Compression Is Geometric, Not Merely Distributional
This compression is not merely a matter of topic frequencies. It is a matter of semantic geometry. Our analysis shows that artist probes occupy a disjoint semantic manifold, with a centroid distance of 2.034 units from the center of public discourse in 8-dimensional space. Even when controlling for style using public probes, this separation persists (1.336 units), and 92.1% of artist probes’ nearest neighbors are other artist probes (p=0.000p=0.000)—far exceeding chance levels and indicating strong neighborhood segregation. This geometric isolation confirms that the mismatch is substantive rather than stylistic. The public is not merely discussing these issues differently; they are talking about different things in different semantic neighborhoods.
4.3 Epistemic Marginalization and the Structure of “Public Noise”
4.3.1 What Occupies the Artist-Absent Regions?
If artists concentrate in 4 clusters, what populates the other 18? Our analysis of artist-absent clusters reveals a discourse landscape dominated by voices and concerns orthogonal to artist experience. This pattern is consistent with epistemic marginalization (Baeyaert, 2025), wherein affected-group knowledge is backgrounded through differential amplification rather than explicit exclusion.
Three discourse types dominate the 14 clusters with zero artist presence:
1. Institutional and market narratives (35.6% of public corpus): These topics feature curators, executives, and institutional voices discussing market implications, funding models, NFT economies, and the philosophical possibilities of AI art, without engaging the labor realities that concern practitioners. This pattern mirrors the actor asymmetries Huang and Gadavanij (2025) document in AI-education discourse, where institutional voices systematically overshadow those of affected practitioners.
2. Technical genealogies and spectacle (18.5% of public corpus): These topics focus on novelty, historical lineage, and technological milestones from DeepDream (Mordvintsev et al., 2015) through GANs (Goodfellow et al., 2014) to the Christie’s Belamy auction. This “origin story” discourse captures public attention through spectacle but does not address consent, compensation, or creative integrity.
3. Legal and policy abstraction (7.3% of public corpus): While governance-relevant, these clusters frame issues through doctrinal, empirical-psychological, or philosophical lenses that exclude the moral and professional framings artists articulate. Legal analysis discusses human authorship thresholds without centering artist testimony; computational creativity philosophy debates machine consciousness without addressing creative labor.
This pattern is consistent with the idea that the public sphere can operate as an exclusionary space, not through active silencing, but through differential amplification. By flooding discourse with technical, institutional, and spectacular narratives, the environment structurally backgrounds practitioners’ experiential knowledge (Huang and Gadavanij, 2025; Zai et al., 2025; Baeyaert, 2025).
4.3.2 Algorithmic Agenda-Setting and the “Post-Mediatized” Sphere
A distinctive feature of the contemporary discourse environment is that agenda-setting is increasingly algorithmic. Sichach (2025) argues that attention is now shaped not only by editorial institutions but also by recommender systems and algorithmic ranking that prioritize engagement and scale. Pane (2025) similarly conceptualizes a “post-mediatized” sphere in which algorithmic curation helps set the epistemic boundaries of what is visible and discussable.
Our results are consistent with—and provide empirical support for—this account. The temporal distribution of public discourse shows that topic salience is driven by a small number of high-visibility events: DeepDream (2015), the Belamy auction (2018), DALL-E release (2021–2022), diffusion model proliferation (2023–2024). These milestone-driven attention spikes generate discourse about capability and spectacle rather than labor impact. Even where artist concerns appear in public discourse (Clusters 9, 17), they are structurally peripheral: Cluster 17 (artist backlash) contains 40.6% of artist probes but only 5.84% of public discourse mass, a $6.95\times$ underrepresentation.
4.4 Why Some Artist Concerns Travel Better Than Others
4.4.1 Threat and Utility Travel Well; Governance-Rights Concerns Do Not
One mechanism for differential salience is narrative portability. Media analyses of AI debates show recurring binary frames (e.g., opportunity versus crisis, tool versus threat) that simplify complexity into repeatable tropes (Bøgh, 2025; Cheung, 2024). These become dominant frames—interpretive packages that achieve salience through repetition and institutional amplification, crowding out alternative viewpoints.
In our data, perceived threat and perceived utility map to Clusters 10 and 16, which together receive 8.69% of public discourse, yielding a salience ratio of $4.81\times$ (artists overrepresent these topics relative to the public). These affective themes—“AI will take my job” or “AI helps my workflow”—translate readily into the binary structure that media discourse favors.
By contrast, governance-rights themes (transparency, ownership, compensation) map predominantly to Cluster 17, which receives only 5.84% of public discourse mass despite containing 40.6% of artist probes, a salience ratio of $6.95\times$. Governance-rights themes require technical pipeline details (e.g., how training data is sourced), legal doctrines (e.g., copyright, licensing, consent), and institutional enforcement mechanisms (e.g., opt-out registries, compensation schemes). These elements are more complex to compress into fast-moving media frames and thus more vulnerable to the compression we document (Sichach, 2025; Zai et al., 2025; Pane, 2025).
This differential marginalization has governance consequences. The concerns most relevant to AI policy—transparency requirements, ownership claims, and compensation mechanisms—are precisely those most underrepresented in public discourse. If policymakers rely on public discourse as a proxy for stakeholder priorities, they will systematically underweight the most actionable demands from artists.
4.4.2 Coalition Mobilization Accentuates the Stakes
Recent organizing by creator coalitions underscores that transparency, consent, and compensation are not fringe concerns but actively contested governance claims. Over 100,000 visual artists in the UK have mobilized through the AOI Coalition against unauthorized use of copyrighted works in AI training (AOI Coalition of Artists, 2025); the US Creators Coalition on AI, launched in December 2025, pursues similar objectives (Creators Coalition on AI, 2025). The Andersen v. Stability AI class action (Andersen et al. v. Stability AI et al., 2023) demonstrates that artist grievances have reached federal litigation, with plaintiffs alleging that generative models trained without consent constitute mass infringement.
These developments strengthen the normative and policy relevance of auditing representational patterns. Artist concerns are not hypothetical or speculative—they are the subject of organized mobilization, legislative testimony, and active litigation. The gap between what artists articulate and what public discourse makes salient is not simply an academic observation; it is a governance failure with material consequences.
4.5 Implications for Policy and Governance
4.5.1 The Risk of Governing for the “Map” Rather Than the “Territory”
The semantic compression we document poses a direct risk to AI governance. If policymakers rely on public sentiment analysis, media monitoring, or “public comment” solicitation to gauge the impact of generative AI on creative labor, they will receive a distorted signal. They will see generic debates about “AI and art” but will miss critical distinctions:
  • Between artists who want opt-out mechanisms and those who want licensing arrangements
  • Between those who would donate freely and those who demand ongoing compensation
  • Between those who accept AI as a tool and those who view it as an existential threat
  • Between those who prioritize transparency and those who prioritize ownership
The 18 ownership frames in our artist sample—ranging from “Artist Exclusive” to “Shared: User + Artist” to “Anti-Ownership/Commons”—all collapse into Cluster 17 in public discourse. Policymakers reading public discourse would not know these distinctions exist.
4.5.2 Auditing Stakeholder Representation as a Governance Practice
Our methodology—projecting stakeholder frames into public discourse space and computing salience ratios—provides an empirical auditing tool. Rather than assuming public discourse reflects stakeholder priorities, we provide measurable indicators of mismatch. This approach could be applied to other technology-affected stakeholder communities (gig workers, content moderators, data subjects) to assess representational equity.
4.6 Generalizability Beyond AI-Art
While our analysis focuses on artists and generative AI, the methodology generalizes to other stakeholder-technology domains, and we are actively validating this transferability. The consensus-based semantic projection approach we introduce provides a stable reference geometry that enables reliable cross-corpus comparison—a capability we are currently applying to stakeholder representation in platform labor and medical AI contexts (methods paper in preparation). The core approach—constructing a semantic reference map from public discourse and projecting stakeholder survey responses to quantify representational alignment—could be applied to:
  1. 1.
    Gig workers and platform labor algorithms: Do public debates about ride-sharing or delivery platforms reflect the concerns of drivers?
  2. 2.
    Patients and medical AI: Does discourse about diagnostic AI capture patient concerns about consent and explainability?
  3. 3.
    Students and educational technology: Are student perspectives represented in debates about AI tutoring systems?
  4. 4.
    Content moderators and platform governance: Do platform policy discussions reflect moderator working conditions?
In each case, the question is whether primary stakeholders achieve proportional representation in the discourse shaping policy, or whether institutional, technical, and spectacular narratives dominate. Our salience-ratio methodology provides a quantitative tool for such audits, enabling systematic comparison across domains and over time.
4.7 Limitations
We acknowledge several limitations:
Methodological constraints: Artist probes are structured responses constrained by survey framing (Lovato et al., 2024), not organic discourse. The public corpus is shaped by Google Search retrieval, potentially overrepresenting highly visible sources. Our use of a single embedding model (e5-large-v2) may introduce representational biases.
Inferential constraints: We document patterns consistent with epistemic marginalization but do not directly demonstrate causal mechanisms. Coalition, agenda-setting, and algorithmic amplification claims remain interpretive without evidence from actor-networks or platform traces.
Scope constraints: Our findings are specific to visual artists and image-generating AI. Generalization to other creative fields requires additional empirical work.
4.8 Conclusion
This study provides a measurement framework for auditing stakeholder representation in technology discourse. We move beyond the normative claim that “artists should be heard” to a measurable demonstration that, in the current public sphere, artist concerns are systematically compressed into a narrow region of meaning-space while institutional, technical, and spectacular narratives occupy the broader discursive landscape.
The empirical signatures are stark:
  • 95% of artist frames concentrate in 4 of 22 topics
  • 14 topics contain zero artist perspective
  • The 14 topics lacking any artist perspectives contain 62.4% of the public corpus
  • Geometric separation persists after style control (centroid distance =1.336=1.336)
[…]
Together, these findings constitute an empirical signature of epistemic marginalization (Baeyaert, 2025)—a structural backgrounding of affected-group knowledge that occurs without explicit silencing or malicious intent, through the combined operation of agenda-setting, algorithmic amplification, and milestone-driven spectacle (Sichach, 2025; Pane, 2025; Cheung, 2024).
For AI governance, the implication is clear: relying on public discourse as a proxy for stakeholder priorities risks legislating for a “map” that does not match the “territory.” The concerns most relevant to policy (transparency, ownership, and compensation) are precisely those most underrepresented.
The public sphere has captured the issue of AI and art. It has not captured the reality of artists’ concerns.
5 Supplementary Methods
Figure M1: Analysis workflow. Three-stage pipeline: (1) corpus construction and embedding, (2) consensus UMAP projection and clustering, (3) cross-corpus comparison and salience analysis.
5.1 Consensus UMAP Implementation
Our consensus-based projection methodology builds on recent computational approaches for comparing stakeholder discourse in shared semantic spaces (Elmholdt et al., 2025; Matsui and Ferrara, 2024). Below we detail the implementation.
Figure S1: Consensus UMAP methodology. We generate 31 UMAP projections with different random seeds, compute pairwise distance matrices for each, average these matrices, and fit a final embedding from the consensus distance structure. This approach increases average seed-to-consensus Adjusted Rand Index from 0.56 (naive coordinate averaging) to 0.71 (distance-matrix consensus), yielding a more stable reference geometry.
5.1.1 Seed Pool and Multi-Run Generation
We generated 31 UMAP projections using seeds 0–30, with the following parameters:
  • n_neighbors: 15
  • min_dist: 0.1
  • n_components: 8
  • metric: cosine
5.1.2 Distance-Matrix Consensus
For each seeded UMAP projection, we computed the pairwise Euclidean distance matrix in 8D space. These 31 matrices were element-wise averaged to produce a consensus distance matrix. A final UMAP was fit using this consensus distance structure as the precomputed distance input.
5.1.3 UMAP Hyperparameter Selection
UMAP hyperparameters were selected via a multi-stage grid search. In the first stage, we searched across two parameters while holding dimensionality fixed:
  • n_neighbors: 9 to 178 in steps of $\sim$9, representing 1% to 20% of the corpus size
  • min_dist: 0.001 to 0.4 in evenly spaced steps
  • n_clusters ($k$): 5 to 30, with optimal $k$ selected via the elbow method on silhouette scores
In the second stage, using the optimal n_neighbors and min_dist from stage one, we performed a sweep across embedding dimensionality:
  • n_components: 4 to 51, representing approximately 0.4% to 5% of the original embedding dimensionality
The dimensionality selection procedure retained configurations in the top 99th percentile of trustworthiness scores, then selected from those the configuration in the top 95th percentile of silhouette scores. This two-stage filtering prioritizes local neighborhood preservation (trustworthiness) while ensuring adequate cluster separation (silhouette). The final selected parameters (n_neighbors=15, min_dist=0.1, n_components=8) achieved strong performance across both criteria.
5.2 Projection Head Architecture
The MLP projection head consists of:
  • Input: 1024-dimensional e5-large-v2 embedding
  • Hidden layers: [512, 256, 128] with ReLU activation and dropout (0.2)
  • Output: 8-dimensional consensus UMAP coordinates
  • Training: Adam optimizer, MSE loss, 80/20 train/validation split
5.3 Robustness Checks
5.3.1 Media Type and Year Associations
  • Media type $\times$ cluster: Cramér’s $V=0.453$
  • Year $\times$ cluster: Cramér’s $V=0.539$
  • Year $\times$ cluster within Articles: Cramér’s $V=0.542$
  • Year $\times$ cluster within Audio: Cramér’s $V=0.433$
  • Year $\times$ cluster within Papers: Cramér’s $V=0.710$
Interpretation: Media and time structure correlate with topics, but do not explain away artist–public divergence; year effects persist within media bins.
5.4 Use of Large Language Models
LLMs were used for the following tasks:
  1. 1.
    Anchor generation: GPT-5.1 generated 250 synthetic Likert-style statements using a factorial design (5 themes $\times$ 5 agreement levels $\times$ 10 discourse styles; see Table S4)
  2. 2.
    Validation filtering: Human + LLM pipeline (Ziems et al., 2024) retained on-theme public probes
  3. 3.
    Topic interpretation: GPT-5.1 Pro, Gemini Pro 1.5, and Claude Opus 4.5 independently labeled clusters (Ziems et al., 2024); GPT-5.1 Pro served as consolidation judge
Human oversight was maintained throughout, and multiple models were used to mitigate single-model biases (Ziems et al., 2024).
6 Supplementary Results
Figure S2: Semantic map of all data in two dimensions. Two-dimensional UMAP projection of the public discourse semantic space showing all 22 topic clusters. Topics containing artist perspectives are highlighted. Artist concerns concentrate in a narrow region of the semantic space, while the majority of topics contain no artist perspective.
6.1 Semantic Pathway Analysis
Analysis of per-artist cross-theme patterns reveals:
  • Top 10 semantic pathways account for 89.4% of artists
  • Compression reflects stable patterns, not idiosyncratic scattering
Figure S3: Semantic pathway concentration analysis. Top 10 pathways account for 89.4% of all artists, indicating stable cross-theme patterns rather than idiosyncratic scattering. Cluster 17 appears in all top 10 pathways for Ownership and Transparency themes; Clusters 10 and 16 capture most Threat and Utility responses.
Figure S4: Theme-to-topic mapping heatmap. Distribution of artist probes (n=1,259n=1{,}259) across public discourse topics by thematic dimension. Affective themes (Threat, Utility) primarily map to Clusters 10 and 16; governance themes (Ownership, Transparency) collapse almost entirely into Cluster 17; compensation themes spread across Clusters 1, 9, and 17. Darker shading indicates higher concentration. All 18 ownership frames—from “Artist Exclusive” to “Anti-Ownership/Commons”—compress into a single topic.
6.2 Base Map Stability Validation
Our consensus UMAP methodology was validated by comparing cluster stability across random seeds:
  • Naive coordinate averaging (31 seeds): Mean ARI =0.56=0.56
  • Distance-matrix consensus (31 seeds): Mean ARI =0.71=0.71
  • Relative improvement: $\sim$27%
The projection head achieves validation $R^{2}=0.73$ and $k$-NN neighborhood preservation of 82.4% ($k=72$).
6.3 Additional Statistical Details
Full permutation test results for geometric separation:
  • Public vs. Artist Probes centroid distance: 2.034 ($p<0.0001$)
  • Public Probes vs. Artist Probes centroid distance: 1.336 ($p<0.0001$)
  • $k$-NN same-source rate (artist probes): 92.1% ($p<0.0001$)
  • $k$-NN same-source rate (public probes vs. artist): 76.5% ($p<0.0001$)
6.4 Search Phrases
Table S1 presents the search phrases used to construct the public discourse corpus.
Table S1: Breadth of search phrases and prevalence of relevant sources. Left: search phrases used to build the public discourse corpus. Right: the number of relevant documents found in the first page of Google search results for each phrase.
Search Phrase Relevant Documents in Top 10 Google Results
AI art 4
AI generated art 5
AI art impact on artists 7
Effects of AI image generators on artists’ careers 5
Artists’ responses to generative AI 7
Artists’ concerns about AI art 5
Interviews with artists about AI image generators 8
How AI affects freelance artists 9
AI copyright challenges for artists 12
Artists suing AI companies 10
How artists adapt to AI tools 6
Artists protest AI training data usage 7
Artists’ opinion on DeepDream art 10
AI art exhibition 8
Art incorporating AI 11
Creatives and AI 9
AI and the creative job market 11
Total 134
6.5 Full Topic Inventory
Table S2 presents the complete 22-topic inventory with labels, keywords, and confidence scores from our multi-annotator interpretation process.
Table S2: Complete list of cluster labels, keywords, and final LLM judge’s confidence in interpretations.
Topic Label Keywords Conf.
0 DeepDream-era neural net art Google DeepDream, neural networks, psychedelic/hallucination, style transfer, authorship 9
1 Industry-backed immersive art tech projection mapping, Panasonic, immersive experience, artist–engineer collaboration 8
2 Eco-humanist immersive AI data art Refik Anadol, Large Nature Model, immersive AI installations, data sculpture 8
3 Harold Cohen’s AARON legacy Harold Cohen, AARON, computational art lineage, machine creativity, NFTs/crypto art 7
4 Panels on digital/immersive art economies digital art economy, responsibility/ownership, metaverse/NFTs, institutional partnerships 7
5 Computational creativity philosophy philosophy of art, human meaning, AI novelty, inter-concept space, cellular automata 8
6 Empirical creativity measurement creativity ratings, semantic distance, divergent thinking, DALL-E 3, co-creation 9
7 Engineering/system theory of AI art digital art systems, AI implementation, intelligent calligraphy, IP/piracy 8
8 GenAI “reality check” 3D artists/workflows, anti-hype critique, hallucinations, labor displacement/rights 8
9 DALL-E 2 “democratization” debate DALL-E 2, democratization narrative, creativity amplifier, attribution/credit 8
10 Pop-culture existential AI art debate Detroit: Become Human, “soul”/mortality, authorship, replacement vs tool 8
11 Luxury art/culture journalism Cartier, Sotheby’s/auctions, Frida Kahlo, public domain, copyright expiration 6
12 Generative art legitimacy and history generative art history, definitions/legitimacy, code + control, feminist pioneers 8
13 AI in the art world: market legitimacy art market/auctions, galleries & collectors, legitimacy/value, AI vs human artist 7
14 GAN-era AI art explainer/history GANs, Obvious/Belamy (Christie’s), Robbie Barrat, latent space 8
15 GenAI tools in education/workflows ChatGPT/GPT-4, creative workflows/prompting, education/classroom use 8
16 AI as artistic collaborator Sougwen Chung, machine hallucination, collective memory, embodied robotics 7
17 Artist backlash and ethical gatekeeping AI resistance/backlash, “tool not agent”, authorship, consent/licensing 8
18 Legal analysis of AI art copyright human authorship requirement, copyrightability, court decisions/precedent 9
19 Museum/institutional Web3/NFT takes museums/institutions, NFTs/Web3, blockchain/metaverse, business models 7
20 Political-economy critique of AI attention economy, data ownership/surveillance, data colonialism, digital commons 8
21 Practical AI adoption in creative education art/design education, tool adoption, prompt engineering, Adobe Firefly 8
6.6 Cluster Distribution by Data Source
Table S3 shows the distribution of public discourse chunks, public probes, and artist probes across all 22 topics. Bold values indicate notable concentrations.
Table S3: Cluster distribution by data source.
Topic Public Public Probes Artist Probes
($n=891$) ($n=379$) ($n=1259$)
0 67 (7.52%) 13 (3.43%) 0
1 71 (7.97%) 14 (3.69%) 25 (1.99%)
2 106 (11.90%) 43 (11.35%) 0
3 23 (2.58%) 6 (1.58%) 0
4 27 (3.03%) 4 (1.06%) 0
5 15 (1.68%) 0 0
6 14 (1.57%) 2 (0.53%) 0
7 10 (1.12%) 1 (0.26%) 0
8 62 (6.96%) 20 (5.28%) 31 (2.46%)
9 40 (4.49%) 21 (5.54%) 187 (14.85%)
10 38 (4.65%) 57 (15.04%) 268 (21.29%)
11 12 (1.34%) 1 (0.26%) 0
12 9 (1.01%) 12 (3.17%) 0
13 67 (7.52%) 13 (3.43%) 0
14 56 (6.29%) 24 (6.33%) 0
15 41 (4.60%) 8 (2.11%) 0
16 36 (4.04%) 55 (14.51%) 235 (18.67%)
17 52 (5.84%) 48 (12.66%) 511 (40.59%)
18 36 (4.04%) 9 (2.37%) 2 (0.16%)
19 34 (3.82%) 1 (0.26%) 0
20 42 (4.71%) 11 (2.90%) 0
21 33 (3.70%) 16 (4.22%) 0
6.7 LLM-Generated Likert Anchor Statements
Table S4 presents representative examples from the 250 synthetic Likert-style anchor statements generated by GPT-5.1, organized by theme and agreement level. Each cell in the 5 $\times$ 5 design (theme $\times$ Likert level) was populated with 10 discourse-style variants drawn from genres present in our public corpus (blog opinion, news editorial, artist interview, legal brief, policy report, panel discussion, first-person narrative, union statement, corporate statement, academic article, etc.). The full set of 250 anchor statements is available as likert_thematic_phrases.csv in our code repository.
Table S4: Design matrix and representative examples of the 250 LLM-generated Likert-style anchor statements (5 themes $\times$ 5 agreement levels $\times$ 10 discourse styles = 250). One example per theme–level combination is shown; the full set is available in our code repository.
Theme Level Representative Example (discourse style)
Utility Str. disagree “From my perspective, artificial intelligence art models are ruining the field of art rather than improving it.” (blog opinion)
Utility Disagree “The panel expressed skepticism that artificial intelligence art models truly benefit the field of art.” (panel discussion)
Utility Neutral “The report takes no position on whether artificial intelligence art models are good or bad for art.” (research report)
Utility Agree “The article argues that artificial intelligence art models open exciting new possibilities for artists.” (news opinion)
Utility Str. agree “This manifesto celebrates artificial intelligence art models as a transformative good for creative practice.” (manifesto)
Ownership Str. disagree “In my view, artificial intelligence generated artwork in an artist’s style should not be considered the artist’s property at all.” (legal opinion)
Ownership Disagree “This blog post argues that users should have primary control over artificial intelligence generated images, rather than the original artists.” (blog opinion)
Ownership Neutral “The article maps the legal uncertainties around ownership of artificial intelligence generated images in an artist’s style.” (academic article)
Ownership Agree “The union calls for recognizing artificial intelligence generated artworks in an artist’s style as that artist’s property.” (union statement)
Ownership Str. agree “The manifesto declares that artificial intelligence generated images in an artist’s style belong entirely to that artist.” (manifesto)
Transparency Str. disagree “Our company rejects proposals to mandate detailed transparency about the art and images used to train our artificial intelligence systems.” (corporate statement)
Transparency Disagree “The brief raises concerns about broad requirements to reveal all art and images used to train artificial intelligence systems.” (policy brief)
Transparency Neutral “Survey respondents were split on whether artificial intelligence model creators should be required to disclose training data details.” (survey report)
Transparency Agree “In the interview, the illustrator says that companies should have to tell us which collections they used to train their artificial intelligence models.” (artist interview)
Transparency Str. agree “The report strongly recommends strict disclosure mandates covering all art and images used to train artificial intelligence models.” (policy report)
Threat Str. disagree “The paper contends that artificial intelligence art models complement, rather than threaten, human artistic labor.” (academic article)
Threat Disagree “In this interview, the painter says they are not overly worried about artificial intelligence art systems taking their job.” (artist interview)
Threat Neutral “The news story quotes some artists who fear displacement by artificial intelligence and others who do not.” (news report)
Threat Agree “Survey data show that many art workers view artificial intelligence art tools as a significant career risk.” (survey report)
Threat Str. agree “Our statement declares that artificial intelligence art models are a direct attack on working artists and their jobs.” (union statement)
Compensation Str. disagree “The report concludes that mandatory artist compensation for artificial intelligence training would be inappropriate.” (policy report)
Compensation Disagree “I am not convinced that artists deserve extensive compensation every time their work informs artificial intelligence art models.” (first person)
Compensation Neutral “Our panel debated revenue sharing, lump sum payments, and public funds for artists affected by artificial intelligence art.” (panel discussion)
Compensation Agree “The paper proposes a licensing model in which artists are paid when their art appears in artificial intelligence training sets.” (academic article)
Compensation Str. agree “This manifesto demands strong profit sharing and royalty rights for artists in all artificial intelligence art contexts.” (manifesto)